OpenStudy (anonymous):

A motor with a power of 7.0 × 104 watts lifts a loaded elevator to a height of 20.00 meters for 40 seconds. What is the force exerted by the motor on the elevator? A. 1.4 × 104 newtons B. 7.0 × 104 newtons C. 1.4 × 105 newtons D. 2.1 × 105 newtons E. 3.5 × 105 newtons

OpenStudy (shubhamsrg):

power = work/time = force*distance/time Just substitute values.

OpenStudy (kropot72):

There are two ways to calculate the work done: 1. Work = watt second =\[7\times 10^{4}\times 40\ joules\ ..........(1)\] 2. Work = Newton meters = \[N \times 20\ joules\ ..........(2)\] To find the value of N we equate equations (1) and (2) and solve for N. \[N \times 20=7\times 10^{4}\times 40\] Note: Don't forget to convert your answer to scientific notation.